OpenStudy (anonymous):

The bottom of a large theater screen is 3ft above your eye level and the top of the acreen is 10ft above your eye level. Assume you walk away from the screen (perpendicular to the screen) at a rate of 3ft/s while looking at the screen. What is the change of the viewing angle when you are 30ft from the wall on which the screen hangs assuming the floor is horizontal ?

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (dumbcow):

i had a diagram but it erased it :{ anyway if you draw it out you should notice 2 triangles using the tangent relationship

OpenStudy (dumbcow):

\[\tan A = \frac{3}{x}\] \[\tan B = \frac{10}{x}\] \[\tan \theta = \tan (B-A) = \frac{\tan B - \tan A}{1 + \tan A \tan B}\]

OpenStudy (dumbcow):

\[\tan \theta = \frac{7x}{x^2 +30}\] now use relation \[\frac{d \theta}{dt} = \frac{d \theta}{dx}*\frac{dx}{dt}\] dx/dt is given as 3

OpenStudy (anonymous):

Ok I get that will try and see :)

OpenStudy (dumbcow):

ok :)

OpenStudy (anonymous):

Where u get the formula for tan(b-a)

OpenStudy (anonymous):

?? :)

OpenStudy (anonymous):

Where does the 30,ft come in

OpenStudy (dumbcow):

that is the "x" value you plug it in at the very end for x after you have dtheta/dt

OpenStudy (anonymous):

How do I set up once I have the 7x/x^2+30

OpenStudy (dumbcow):

you have to differentiate.... use quotient rule http://en.wikipedia.org/wiki/Quotient_rule

OpenStudy (perl):

nice problem

OpenStudy (dumbcow):

you should get \[\sec^2 \theta \frac{d \theta}{dx} = \frac{210 - 7x^2}{(x^2 +30)^2}\]

OpenStudy (anonymous):

Ok yes caught up!

OpenStudy (dumbcow):

ok :) now if we solve for d/dx and multiply by 3, we will have rate angle is changing \[\frac{d \theta}{dt} = 3 \cos^2 \theta \frac{210-7x}{(x^2 +30)^2}\] last part is finding cos^2 in terms of x

OpenStudy (dumbcow):

to do this imagine we have a triangle and use the tan relation  \[\cos^2 \theta = \frac{(x^2 +30)^2}{h^2}\] find "h" using pythagorean thm

OpenStudy (anonymous):

85.02 so 85

OpenStudy (dumbcow):

what is 85?

OpenStudy (anonymous):

How we solve for h we know the 30ft not plug in for x

OpenStudy (dumbcow):

oh well no you should prob wait til very end to plug in x=30 get "h" in terms of x something went wrong with your computing, it wouldnt be 85 even if you plug in x=30

OpenStudy (anonymous):

H= sqrt (x^4+109x^2+900) Please lol

OpenStudy (dumbcow):

right

OpenStudy (anonymous):

Thank god lol

OpenStudy (dumbcow):

haha so now you have \[\frac{d \theta}{dt} = 3 (\frac{210 - 7x^2}{x^4 +109 x^2 +900})\] plug in x =30 and you're done

OpenStudy (anonymous):

Should the bottom be still under square root

OpenStudy (dumbcow):

no because remember we wanted cos^2 which has "h^2" on bottom the square cancels out the sqrt

OpenStudy (anonymous):

Right haaha

OpenStudy (anonymous):

U are the best !!